# On the Generalization vs Fidelity Paradox in Knowledge Distillation

### v1

@@ Abstract
Knowledge distillation (KD) is a key technique for compressing large language models into smaller ones while preserving performance. Despite the recent traction of KD research, its effectiveness for smaller language models (LMs) and the mechanisms driving knowledge transfer remain underexplored. In this work, we present the first large-scale empirical and statistical analysis of KD across models ranging from 0.5B to 7B parameters on 14 complex reasoning tasks in a zero-shot setting. Our findings reveal that KD can improve the average performance of smaller models by up to , with a peak task specific gain of , while providing only marginal benefits () for larger models. Surprisingly, teacher performance has a minimal impact on student outcomes, while teacher task expertise impacts KD effectiveness. A correlation study indicates that smaller LMs benefit more from KD, whereas larger LMs show diminished gains. Additionally, we uncover a misalignment between improvements in student performance and reasoning fidelity, suggesting that while KD enhances accuracy, it does not always maintain the structured decision-making processes of the teacher. Our ablation study further highlights the importance of teacher signals and logit smoothing in influencing students’ performance after distillation. Overall, our study offers a comprehensive empirical and statistical assessment of KD, highlighting both its benefits and trade-offs when distilling knowledge from larger to smaller LMs.
@@ 1 Introduction
In this paper, we conduct a large-scale empirical and statistical analysis to evaluate the effectiveness of knowledge distillation for small LMs ranging from 0.5B to 7B parameters across complex reasoning tasks in a zero-shot setting. We investigate the impact of KD on the generalizability of small LMs to explore different aspects of knowledge distillation. We summarize the key findings from our study as follows111The source code of our study is made available at https://github.com/LCS2-IIITD/KD_generalization.:
@@ 2 Related Work
KD has been widely used to compress LLMs while preserving performance Hinton et al. (2015). Traditional KD relies on soft labels to transfer knowledge, with extensions such as feature-based Romero et al. (2014), self-distillation Furlanello et al. (2018), and multi-teacher distillation You et al. (2017). However, these approaches struggle with autoregressive sequence generation due to exposure bias and knowledge loss in smaller models. To address this, Generalized Knowledge Distillation trains students on self-generated sequences, mitigating distribution mismatch Agarwal et al. (2024). Instruction-tuning-CoT enables smaller models to inherit reasoning capabilities via instruction-based fine-tuning Ranaldi and Freitas (2024). Multi-teacher KD aggregates strategies from multiple models, improving generalization Tian et al. (2024), while MiniLLM prioritizes high-probability teacher outputs via reverse KL divergence, reducing overfitting Gu et al. (2024). Adaptive Teaching KD (ATKD) adjusts knowledge transfer dynamically based on token difficulty, preventing degradation in large teacher-student setups Zhong et al. (2024). Despite advancements, most KD methods prioritize performance gains over understanding why certain strategies succeed or fail. Gaining deeper insights into KD effectiveness is crucial for developing more reliable, generalizable models.
@@
Unlike prior works that primarily focus on task-specific knowledge distillation or performance benchmarking on a limited set of tasks, this study provides a large-scale empirical and statistical analysis of KD across diverse reasoning benchmarks with small LMs of varying sizes and capabilities. While most existing studies overlook the explainability of KD, our work systematically examines the key factors influencing its effectiveness, providing valuable practical insights into the process.
@@ 3 Methodology
4 Experimental Setup
In our empirical study, we use Qwen-2.5 Yang et al. (2024) (0.5B, 1.5B, 3B, 7B, 14B) and LLaMA-3 Dubey et al. (2024) (3.2-1B, 3.2-3B, 3.1-8B) model series, with all pretrained weights sourced from Huggingface Wolf et al. (2020), evaluated on mathematical and commonsense reasoning tasks. For mathematical reasoning, we fine-tune base models on Math10K Hu et al. (2023) and evaluate on GSM8K Cobbe et al. (2021), SVAMP Patel et al. (2021), MultiArith Roy and Roth (2015), AddSub Hosseini et al. (2014), AQuA Ling et al. (2017), and SingleEq Koncel-Kedziorski et al. (2015). For commonsense reasoning, we fine-tune on Commonsense15K Hu et al. (2023) and evaluate on Hellaswag Zellers et al. (2019), Winogrande Sakaguchi et al. (2021), ARC Clark et al. (2018), OBQA Mihaylov et al. (2018), BoolQ Clark et al. (2019), PiQA Bisk et al. (2020), and SiQA Sap et al. (2019). Detailed dataset descriptions and splits are provided in Appendix A and Table 6. We use LoRA Hu et al. (2021) adapters with and for supervised fine-tuning and KD fine-tuning of all the models. All the experiments were conducted on a single Nvidia-A100 GPU.
@@ 4 Experimental Setup
[…]
@@
In this paper, along with the performance measurement of the student models, we use teacher-student agreement and reasoning fidelity scores to understand the alignment between teacher and student post-KD. Agreement quantifies how often the student replicates the teacher’s outputs and is measured using top-1 agreement (fraction of matching predictions). However, this metric only relies on the final answer and does not measure the quality of the intermediate reasoning steps. Reasoning fidelity, on the other hand, captures how well the student mirrors the teacher’s reasoning process rather than just final predictions. We use BLEU Papineni et al. (2002) score between teacher and student reasoning outputs to compute the fidelity.
@@ 5 Experimental Results
Figure 1 shows that KD consistently outperforms SFT in both mathematical and commonsense reasoning tasks, particularly benefiting smaller models. For Qwen-0.5B, KD improves performance by and for math and commonsense reasoning respectively, but this effect diminishes as student size increases (, for Qwen-1.5B, , for Qwen-3B and , for Qwen-7B). Similar performance improvements are observed with LLaMA student models – and for LLaMA-1B and and for LLaMA-3B model. Among KD methods, RevKD is the most consistent, yielding an average gain of across all student sizes. The statistical t-test results (Table 1) confirm these improvements as significant (-value ). However, a one-way ANOVA test shows no significant difference between KD methods, indicating that all KD methods are equally effective.
@@
Table 2 reports the Spearman rank correlation between student improvement after KD and teacher performance. In mathematical reasoning, structured tasks like AddSub (0.30, -value=) and SingleEq (0.35, -value=) show positive correlations, suggesting stronger teachers enhance student performance. However, complex reasoning tasks such as GSM8k (, -value=) and AQuA (, -value=) exhibit weak or negative correlations, indicating that teacher quality alone does not dictate KD success. In commonsense reasoning, correlations are generally weak or negative, implying minimal influence of teacher performance. Overall, teacher quality inconsistently impacts student gains, with structured mathematical tasks benefiting more than open-ended reasoning tasks. However, Figure 2 shows that a task-unaware teacher can significantly degrade student performance. The post-distillation performance of the student model can drop up to , if the teacher is not fine-tuned on downstream domain.
@@
Table 3 shows a strong negative correlation (, -value=) between KD improvement and student size, indicating diminishing returns as model size increases. This trend is most pronounced in MultiArith (, -value=), AddSub (, -value=) and SingleEq (, -value=), where larger models already possess strong reasoning capabilities. In commonsense reasoning, similar effect is observed (, -value=). These results confirm that KD benefits smaller LMs significantly, while larger LMs see diminishing returns as they already exhibit strong reasoning.
@@
How does KD impair teacher-student agreement?
Figure 3 and Figure 8 of Appendix B show that post-KD agreement between student and teacher models declines as student size increases. Smaller models (e.g., Qwen-0.5B and Qwen-1.5B) exhibit higher agreement with larger Qwen-14B model in structured mathematical tasks like AddSub () and MultiArith (), indicating effective transfer of well-defined rules. Table 4a reveals that higher agreement does not always translate to better student performance (e.g., negative correlation for MultiArith and SVAMP) suggesting that effective students deviate from teacher outputs rather than mimicking them. For commonsense reasoning (reported in Table 4b), the correlation between teacher-student agreement and the student performance remains weak, indicating statistically insignificant correlation.
@@
Teacher-student fidelity, measuring alignment in reasoning patterns, varies by model size and distillation method. While KD improves fidelity over SFT (Figure 4), Table 5 indicates weak correlations between fidelity and student performance, implying that strict imitation does not necessarily enhance performance. For instance, AddSub (, -value=) and GSM8k (, -value=) show low correlations, reinforcing that high fidelity alone is not a reliable predictor of KD effectiveness. Error analysis in Table 9 of Appendix C further highlights the feeble connection between teacher-student fidelity and student generalization.
@@ Appendix A Datasets
The train/val/test dataset splits for mathematical and commonsense reasoning datasets are highlighted in Table 6a and 6b, respectively.
@@ Appendix B Results
We report the detailed results for SFT, SeqKD, RevKD and GKD of all teacher-student combinations for Math Reasoning tasks in Table 7 and Commonsense reasoning tasks in Table 8. RevKD consistently delivers the best performance, with Qwen-3B distilled from Qwen-7B achieving , outperforming SeqKD () and GKD (). However, larger student models benefit less from KD, as seen in Qwen-7B distilled from Qwen-14B, which only shows a marginal improvement over its fine-tuned counterpart ( ). Structured tasks such as MultiArith and SingleEq exhibit the highest gains, often exceeding accuracy, indicating that KD effectively transfers arithmetic-based reasoning. Conversely, AQuA and GSM8K remain more challenging, particularly for smaller students, highlighting that complex multi-step reasoning is harder to distill. Additionally, the results suggest that KD effectiveness does not depend on model architecture, as LLaMA-based students benefit similarly from KD, with LLaMA-1B distilled from LLaMA-8B with GKD achieving imporving the SFT model by . Finally, distillation from stronger teachers improves student performance, but not linearly, as Qwen-3B distilled from Qwen-7B performs almost as well as when distilled from Qwen-14B, reinforcing that teacher expertise matters more than sheer size. Overall, these findings underscore that KD significantly enhances smaller models’ reasoning capabilities, but diminishing returns appear for larger models, necessitating adaptive KD strategies tailored to student capacity and task complexity.
@@
Figure 9 and 11 illustrates the teacher-student fidelity for Qwen-14B and LLaMA-8B teacher models, respectively. The reasoning fidelity analysis highlights how closely student models replicate their teacher’s reasoning process post-distillation. Across all models, KD improves fidelity over SFT, demonstrating that knowledge transfer enhances reasoning similarity. With Qwen2.5-0.5B model distilled from Qwen2.5-14B, fidelity in GSM8K increases from (SFT) to (KD), and in SingleEq, it improves from to . Similarly, for the larger Qwen2.5-3B model, fidelity scores increase across tasks, such as MultiArith (from to ) and SVAMP (from to ). However, the improvements are modest for larger students, suggesting that they develop their own reasoning strategies instead of strictly mimicking the teacher.

### v2

@@ Abstract
Knowledge distillation (KD) is a key technique for compressing large language models into smaller ones while preserving performance. Despite the recent traction of KD research, its effectiveness for smaller language models (LMs) and the mechanisms driving knowledge transfer remain underexplored. In this work, we present the first large-scale empirical and statistical analysis of KD across models ranging from 0.5B to 7B parameters on 19 complex reasoning and instruction following tasks in a zero-shot setting. Our findings reveal that KD can improve the average performance of smaller models by up to , with a peak task specific gain of , while providing only marginal benefits () for larger models. Surprisingly, teacher performance has a minimal impact on student outcomes, while teacher task expertise impacts KD effectiveness. A correlation study indicates that smaller LMs benefit more from KD, whereas larger LMs show diminished gains. Additionally, we uncover a misalignment between improvements in student performance and reasoning fidelity, suggesting that while KD enhances accuracy, it does not always maintain the structured decision-making processes of the teacher. Our ablation study further highlights the importance of teacher signals and logit smoothing in influencing students’ performance after distillation. Overall, our study offers a comprehensive empirical and statistical assessment of KD, highlighting both its benefits and trade-offs when distilling knowledge from larger to smaller LMs.
@@ 1 Introduction
In this paper, we conduct a large-scale empirical and statistical analysis to evaluate the effectiveness of knowledge distillation for small LMs ranging from 0.5B to 7B parameters across complex reasoning and instruction following tasks in a zero-shot setting. We investigate the impact of KD on the generalizability of small LMs to explore different aspects of knowledge distillation. We summarize the key findings from our study as follows111The source code of our study is made available at https://github.com/LCS2-IIITD/KD_generalization.:
@@ 2 Related Work
KD has been widely used to compress LLMs while preserving performance Hinton et al. (2015). Traditional KD relies on soft labels to transfer knowledge, with extensions such as feature-based Romero et al. (2014), self-distillation Furlanello et al. (2018), multi-teacher distillation You et al. (2017) and collaborative distillation (Sengupta et al., 2023). However, these approaches struggle with autoregressive sequence generation due to exposure bias and knowledge loss in smaller models. To address this, Generalized Knowledge Distillation trains students on self-generated sequences, mitigating distribution mismatch Agarwal et al. (2024). Instruction-tuning-CoT enables smaller models to inherit reasoning capabilities via instruction-based fine-tuning Ranaldi and Freitas (2024). Multi-teacher KD aggregates strategies from multiple models, improving generalization Tian et al. (2024), while MiniLLM prioritizes high-probability teacher outputs via reverse KL divergence, reducing overfitting Gu et al. (2024). Adaptive Teaching KD (ATKD) adjusts knowledge transfer dynamically based on token difficulty, preventing degradation in large teacher-student setups Zhong et al. (2024). Despite advancements, most KD methods prioritize performance gains over understanding why certain strategies succeed or fail. Gaining deeper insights into KD effectiveness is crucial for developing more reliable, generalizable models.
@@
Unlike prior works that primarily focus on task-specific knowledge distillation or performance benchmarking on a limited set of tasks, this study provides a large-scale empirical and statistical analysis of KD across diverse reasoning benchmarks with small LMs of varying sizes and capabilities. While most existing studies overlook explainability of KD, our work systematically examines the key factors influencing its effectiveness, providing valuable practical insights into the process.
@@ 3 Methodology
| ANOVA all methods | 0.49 (0.6) | 0.72 (0.5) | 0.20 (0.8) | 0.18 (0.8) | 0.04 (1.0) | 0.45 (0.6) | 1.12 (0.3) | 0.27 (0.8) | 0.30 (0.8) |
| Test Type | Dolly | Self | SNI | UNI | Vicuna | Average |
| t-test all methods | 6.31 (0.0) | 4.85 (0.0) | 7.01 (0.0) | 8.23 (0.0) | 5.35 (0.0) | 7.18 (0.0) |
| t-test GKD | 14.09 (0.0) | 10.73 (0.0) | 11.25 (0.0) | 18.2 (0.0) | 11.74 (0.0) | 25.66 (0.0) |
| t-test RevKD | 16.63 (0.0) | 8.21 (0.0) | 15.69 (0.0) | 35.42 (0.0) | 12.17 (0.0) | 34.02 (0.0) |
| t-test SeqKD | -2.63 (1.0) | -1.85 (0.9) | -0.32 (0.6) | 1.55 (0.1) | -3.63 (1.0) | -1.66 (0.9) |
| ANOVA all methods | 117.82 (0.0) | 40.05 (0.0) | 67.28 (0.0) | 196.88 (0.0) | 85.16 (0.0) | 329.47 (0.0) |
4 Experimental Setup
In our empirical study, we use Qwen-2.5 Yang et al. (2024) (0.5B, 1.5B, 3B, 7B, 14B) and LLaMA-3 Dubey et al. (2024) (3.2-1B, 3.2-3B, 3.1-8B) model series, with all pretrained weights sourced from Huggingface Wolf et al. (2020), evaluated on mathematical reasoning, commonsense reasoning and instruction following tasks. For mathematical reasoning, we fine-tune base models on Math10K Hu et al. (2023) and evaluate on GSM8K Cobbe et al. (2021), SVAMP Patel et al. (2021), MultiArith Roy and Roth (2015), AddSub Hosseini et al. (2014), AQuA Ling et al. (2017), and SingleEq Koncel-Kedziorski et al. (2015). For commonsense reasoning, we fine-tune on Commonsense15K Hu et al. (2023) and evaluate on Hellaswag Zellers et al. (2019), Winogrande Sakaguchi et al. (2021), ARC Clark et al. (2018), OBQA Mihaylov et al. (2018), BoolQ Clark et al. (2019), PiQA Bisk et al. (2020), and SiQA Sap et al. (2019). For instruction following, we fine-tune using Dolly-15K Gu et al. (2024) and evaluate on Dolly, SelfInst Wang et al. (2022a) (denoted by ‘self’ throughout the paper), Vicuna Chiang et al. (2023), SNI Wang et al. (2022b) and UNI Honovich et al. (2023). Detailed dataset descriptions and splits are provided in Appendix A and Table 6. We use LoRA Hu et al. (2021) adapters with and for supervised fine-tuning and KD fine-tuning of all the models. All the experiments were conducted on a single Nvidia-A100 GPU.
@@ 4 Experimental Setup
| Test Type | Dolly | Self | SNI | UNI | Vicuna | Average |
| Spearman rank | 0.16 (0.4) | 0.04 (0.8) | 0.16 (0.4) | 0.03 (0.9) | -0.0 (1.0) | 0.14 (0.4) |
@@
In this paper, along with the performance measurement of student models, we use teacher-student agreement and reasoning fidelity scores to understand alignment between teacher and student post-KD. Agreement quantifies how often student replicates the teacher’s outputs and is measured using top-1 agreement (fraction of matching predictions). However, this metric only relies on the final answer and does not measure the quality of the intermediate reasoning steps. Reasoning fidelity, on the other hand, captures how well the student mirrors the teacher’s reasoning process rather than just final predictions. We use BLEU Papineni et al. (2002) score between teacher and student reasoning outputs to compute fidelity.
@@ 5 Experimental Results
| Test Type | Dolly | Self | SNI | UNI | Vicuna | Average |
| Spearman rank | 0.08 (0.7) | 0.3 (0.1) | -0.32 (0.1) | -0.02 (0.9) | 0.12 (0.5) | -0.03 (0.9) |
@@
Figure 1 shows that KD consistently outperforms SFT in all three benchmarks - mathematical reasoning, commonsense reasoning and instruction following , particularly benefiting smaller models. For Qwen-0.5B, KD improves performance by and for math and commonsense reasoning respectively and absolute points for instruction following, but this effect diminishes as student size increases (, , for Qwen-1.5B, , , for Qwen-3B and , , for Qwen-7B). Similar performance improvements are observed with LLaMA student models – , and for LLaMA-1B and , and for LLaMA-3B model. Among KD methods, RevKD is the most consistent, yielding an average gain of across all student sizes. The statistical t-test results (Table 1) confirm these improvements as significant (-value ). While one-way ANOVA test reveals a difference between KD methods on instruction-following tasks, it shows no significant differences on mathematical and commonsense reasoning tasks—suggesting that all KD methods perform similarly on these two benchmarks.
@@
Table 2 reports the Spearman rank correlation between student improvement after KD and teacher performance. In mathematical reasoning, structured tasks like AddSub (0.30, -value=) and SingleEq (0.35, -value=) show positive correlations, suggesting stronger teachers enhance student performance. However, complex reasoning tasks such as GSM8k (, -value=) and AQuA (, -value=) exhibit weak or negative correlations, indicating that teacher quality alone does not dictate KD success. In commonsense reasoning and instruction following tasks, correlations are generally weak or negative, implying minimal influence of teacher performance. Overall, teacher quality inconsistently impacts student gains, with structured mathematical tasks benefiting more than open-ended reasoning tasks. However, Figure 2 shows that a task-unaware teacher can significantly degrade student performance. Post-distillation performance of student model can drop up to , if teacher is not fine-tuned on downstream domain.
@@
Table 3 shows a strong negative correlation (, -value=) between KD improvement and student size, indicating diminishing returns as model size increases. This trend is most pronounced in MultiArith (, -value=), AddSub (, -value=) and SingleEq (, -value=), where larger models already possess strong reasoning capabilities. In commonsense reasoning, similar effect is observed (, -value=). In instruction following, correlations are mostly weak. These results confirm that KD benefits smaller LMs significantly, while larger LMs see diminishing returns as they already exhibit strong reasoning.
How does KD impair teacher-student agreement?
Figure 3 and Figure 8 of Appendix B show that post-KD agreement between student and teacher models declines as student size increases. Smaller models (e.g., Qwen-0.5B and Qwen-1.5B) exhibit higher agreement with larger Qwen-14B model in structured mathematical tasks like AddSub () and MultiArith (), indicating effective transfer of well-defined rules.
@@
| Test Type | Dolly | Self | SNI | UNI | Vicuna |
| Spearman rank | 0.3 (0.2) | -0.06 (0.8) | -0.06 (0.8) | -0.13 (0.6) | 0.07 (0.8) |
Table 4a reveals that higher agreement does not always translate to better student performance (e.g., negative correlation for MultiArith and SVAMP) suggesting that effective students deviate from teacher outputs rather than mimicking them. For commonsense reasoning (reported in Table 4b), the correlation between teacher-student agreement and the student performance remains weak, indicating statistically insignificant correlation.
@@
Teacher-student fidelity, measuring alignment in reasoning patterns, varies by model size and distillation method. Along with the BLEU-based fidelity metric, we devise an alternative fidelity metric, calculated using the cosine similarity between sentence embeddings (Reimers and Gurevych, 2019) of teacher and student responses. The Spearman correlation between BLEU scores and cosine similarity is 0.97 (p-value = 1e-9), indicating strong agreement. Therefore, we report only BLEU-based fidelity metric, as the resulting hypotheses remain unchanged. While KD improves fidelity over SFT (Figure 4a), Table 5 indicates weak correlations between fidelity and student performance, implying that strict imitation does not necessarily enhance performance. For instance, AddSub (, -value=) and GSM8k (, -value=) show low correlations, reinforcing that high fidelity alone is not a reliable predictor of KD effectiveness. Error analysis in Table 10 of Appendix C further highlights the feeble connection between teacher-student fidelity and student generalization.
@@
Impact of teacher-student gap on KD.
Mirzadeh et al. (2020) observe that in computer vision models, student performance tends to degrade when the gap between the teacher and student is too large. We find a similar trend in our setting: the most effective teacher for a student model is not necessarily the largest one available. For instance, as shown in Table 7, Qwen-1.5B achieves its best performance when distilled from Qwen-7B rather than the larger Qwen-14B. We observe this pattern across multiple knowledge distillation methods and datasets.
@@ 6 Conclusion
Acknowledgment
We acknowledge the support of the IBM-IITD AI Horizons network. T. Chakraborty acknowledges the support of the Rajiv Khemani Young Faculty Chair Professorship in Artificial Intelligence.
@@ Appendix A Datasets
The Instruction following benchmark consists of the following datasets:
Self Wang et al. (2022a) - This dataset consists of 252 user-oriented instruction-following samples.
Vicuna Chiang et al. (2023) - This dataset contains 80 challenging questions synthesized by GPT-4 used in Vicuna evaluation.
SNI Wang et al. (2022b) - The dataset comprises approximately 9K samples drawn from around 119 tasks within the Super-Natural Instructions benchmark. Following Gu et al. (2024), we divide the samples into three subsets based on the length of the ground-truth responses. For our experiments, we use the subset with response lengths in the range [11, ] which contains about 1.6K samples.
UNI Honovich et al. (2023) - This dataset consists of samples from the core set of Unnatural Instructions. As with S-NI, we focus on the subset where ground-truth response lengths fall within the range [11, ]. We use first 2k samples of them for evaluation.
The train/val/test dataset splits for mathematical, commonsense reasoning and instruction following datasets are highlighted in Table 6.
@@
| Dataset | # train | # validation | # test |
| Dolly15K | 12.5K | 1k | 500 |
| SelfInst | - | - | 252 |
| VicunaEval | - | - | 80 |
| S-NI | - | - | 1.6K |
| UnNI | - | - | 2k |
@@ Appendix B Results
We report the detailed results for SFT, SeqKD, RevKD and GKD of all teacher-student combinations for Math Reasoning tasks in Table 7, Commonsense reasoning tasks in Table 8 and Instruction following tasks in Table 9. RevKD consistently delivers the best performance, with Qwen-3B distilled from Qwen-7B achieving , outperforming SeqKD () and GKD (). However, larger student models benefit less from KD, as seen in Qwen-7B distilled from Qwen-14B, which only shows a marginal improvement over its fine-tuned counterpart ( ). Structured tasks such as MultiArith and SingleEq exhibit the highest gains, often exceeding accuracy, indicating that KD effectively transfers arithmetic-based reasoning. Conversely, AQuA and GSM8K remain more challenging, particularly for smaller students, highlighting that complex multi-step reasoning is harder to distill. Additionally, the results suggest that KD effectiveness does not depend on model architecture, as LLaMA-based students benefit similarly from KD, with LLaMA-1B distilled from LLaMA-8B with GKD achieving imporving the SFT model by . Finally, distillation from stronger teachers improves student performance, but not linearly, as Qwen-3B distilled from Qwen-7B performs almost as well as when distilled from Qwen-14B, reinforcing that teacher expertise matters more than sheer size. Overall, these findings underscore that KD significantly enhances smaller models’ reasoning capabilities, but diminishing returns appear for larger models, necessitating adaptive KD strategies tailored to student capacity and task complexity.
@@
On instruction following tasks as well, RevKD outperforms other KD methods. Distilled Qwen-0.5B improves highest on SNI and UNI by 0.11 and 0.15 absolute points respectively. Similarly other distilled Qwen and Llama models achieve largest improvements on these two tasks.
@@
Figure 9 and 11 illustrates the teacher-student fidelity for Qwen-14B and LLaMA-8B teacher models, respectively. The reasoning fidelity analysis highlights how closely student models replicate their teacher’s reasoning process post-distillation. Across all models, KD improves fidelity over SFT, demonstrating that knowledge transfer enhances reasoning similarity. With Qwen2.5-0.5B model distilled from Qwen2.5-14B, fidelity in GSM8K increases from (SFT) to (KD), and in SingleEq, it improves from to . Similarly, for the larger Qwen2.5-3B model, fidelity scores increase across tasks, such as MultiArith (from to ) and SVAMP (from to ). However, the improvements are modest for larger students, suggesting that they develop their own reasoning strategies instead of strictly mimicking the teacher. Similarly as seen in Figure 9b and Figure 11b, KD significantly improves fidelity over SFT across all instruction following tasks in all the models. SNI and UNI show the largest improvement among instruction following tasks.